OpenStudy (toxicsugar22):

Andy drives 300 miles to Lake Tahoe at 70 miles per hour and returns home at 50 mph. What is his average speed for the round trip. (It is not 6o miles per hour.) Write expressions for the time it takes for each leg of the trip if Andy drives a distance d, at speed r base 1 and returns at speed r base 2

OpenStudy (anonymous):

300/70 is the total time to get out there. 300/50 is the total time to get back. In variables, distance = 300 speed out = r1 speed back = r2 so you have [(d/r1)+(d/r2)]/2

OpenStudy (anonymous):

er, I don't think that's right actually...d/r1 gives the amount of time it takes to go out, d/r2 gives the amount of time it takes to come back. To find the average, you need the total distance traveled (d*2) divided by the total time it took (d/21+d/r2)

OpenStudy (toxicsugar22):

I dont think the total time is right

OpenStudy (toxicsugar22):

r u sure its right

OpenStudy (toxicsugar22):

u sure

OpenStudy (anonymous):

300mi/70mph = 4.286 hours to get there 300mi/50mph = 6.00 hours to get back Total time traveled = 4.286 + 6.000 = 10.286 hours Total distance traveled = 300*2 = 600 mi Average speed = distance/time = 600/10.286 = 58.33mph